In 2018, the braking distance of Toyota Camry cars on a wet surface follows a normal distribution. Its mean is 122 feet with a standard deviation of 20 feet. What is the probability that a randomly selected Toyota Camry will have a braking distance of more than 130 feet?

1
Expert's answer
2022-02-02T09:36:08-0500

Let "X=" the braking distance:"X\sim N(\mu, \sigma^2)."

Given "\mu=122ft, \sigma=20ft."


"P(X>130)=1-P(X\le130)"

"=1-P(Z\leq\dfrac{130-122}{20})=1-P(Z\leq0.4)"

"\approx0.344578"